Find the volume generated by revolving the area enclosed by the curve x2 + y2 = 9 about

the x-axis.


1
Expert's answer
2021-12-23T12:31:01-0500

Curve: "x^{2}+y^{2}=9"

It is an equation of the circle, with radius 3 and centered at the origin.

If you notate a circle about x-axis,


Yow will get a sphere of the same radius and also centered at the origin.

so, volume: "V=\frac{4}{3} \pi r^{3}" (volume of sphere)

Β "\begin{aligned}

\Rightarrow V &=\frac{4}{3} \pi(3)^{3} \quad[\because \text { radius }=3] \\

&=36 \pi \\

\therefore \quad \text {Volume } &=36 \pi 

\end{aligned}"